• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name X. Zheng

    Log entry time 16:03:04 on April 01, 2010

    Entry number 315517

    keyword=Day shift summary 4/1

    08:00 shift started with psudo-production 50uA on empty target. Beam position is (0,1) at A and (1,2) at B;

    08:20 MCC called and wanted to send us 100uA. We verified the empty target, consulted the RC and said yes. THe first run with 100uA is 2483;

    09:00 Beam study started. We asked for controlled access with full survey.

    10:20 Full survey is finished and the hall is in controlled access;

    13:38 Everyone is out of the Hall in preparation for invasive beam study (invasive to Hall A).

    16:00 Shift ended with beam study, no beam in Hall A.

    Run Summary (pseudo-production):

    2482empty50132kended early to increase beam current
    2485empty50110kended early for beam study

    Other runs during this shift are either junk or pedestal runs.